OpenStudy (milliedelongg):

Need help with a range of problems for my college algebra class. I understand none of it and need helo asap.

OpenStudy (asc.bchs):

I might be able to help. What question?

OpenStudy (milliedelongg):

There are so many, but i will show you the first one.

OpenStudy (milliedelongg):

x+5y-z=-8 x+6y+z=-19 x-7y+6z=0 (And they have a big { on the right side idk if it matters or not)

OpenStudy (milliedelongg):

Another problem is 3x+4y-5z=8 4x+y-z=-1 x-3y+3z=-10

OpenStudy (asc.bchs):

What are the directions?

OpenStudy (milliedelongg):

Use right-to-left elimination to solve the system of equations